OpenStudy (anonymous):

Billy has 1 gallon of paint. He is going to pour it into a paint tray that measures 10 inches wide, 14 inches long, and 4 cm deep.

hartnn (hartnn):

what do you need help with ?

OpenStudy (anonymous):

Billy has 1 gallon of paint. He is going to pour it into a paint tray that measures 10 inches wide, 14 inches long, and 4 cm deep. (1 gallon = 231 in3, 1 inch = 2.54 cm) Which of the following scenarios will happen? The paint will not fill the tray by 441 cm3. The paint will not fill the tray by 10.53 in3. The paint will fill the tray exactly. The paint will overfill the tray by 10.53 in3.

OpenStudy (anonymous):

my question has been around for an hour and i still have no replies

OpenStudy (anonymous):

mine too

OpenStudy (anonymous):

i have so many question and no one helps me

OpenStudy (anonymous):

i managed to help out 2 other people

OpenStudy (anonymous):

yeah but i need help now

hartnn (hartnn):

volume of tray = length *base*height = 10*14*4 = ... ?

OpenStudy (anonymous):

see i dont know

OpenStudy (anonymous):

im not good at math

hartnn (hartnn):

wait, first you need to convert the 4cm into inches so, if 1 inch =2.54 cm, 4cm = ...? inch

OpenStudy (anonymous):

use a calculator

OpenStudy (anonymous):

im not good at math i thought you guys would give me an answer because i have a max time limit and i dont have one

hartnn (hartnn):

sorry, we do not give answers this is not an answering site, this is a learning site.

OpenStudy (anonymous):

just google or search for an online calculator

OpenStudy (anonymous):

well everyone else is giving answers i just need one and thanks ill do that

hartnn (hartnn):

read this : http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

lol sorry i dont really care i just need help finding the answer

OpenStudy (anonymous):

once you find an online calculator 1. convert 1 inch =2.54 cm, 4cm = ...? inch 2.find volume of tray = length *base*height = 10*14*4 = ...

OpenStudy (anonymous):

how would i convert

OpenStudy (anonymous):

4 * 2.54

OpenStudy (anonymous):

10.16?

hartnn (hartnn):

4/2.54 actually

OpenStudy (anonymous):

so i divide?

OpenStudy (anonymous):

yeah sorry about that

OpenStudy (anonymous):

1.57

OpenStudy (anonymous):

now find the volume of the tray

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

10*14*4

OpenStudy (anonymous):

560

hartnn (hartnn):

actually, 10*14*1.57

OpenStudy (anonymous):

219.8?

hartnn (hartnn):

yes. 219.8 cm^3 but 1 gallon is 231, cm^3 so, how much paint will overfill ?

OpenStudy (anonymous):

441?

hartnn (hartnn):

what is 231-219.8 = ..?

OpenStudy (anonymous):

11.2

hartnn (hartnn):

and the nearest choice to 11.2in^3 is The paint will overfill the tray by 10.53 in3.

OpenStudy (anonymous):

so it will overfill correct

hartnn (hartnn):

yup , because 1 gallon is 231, but the vesel is just of 219...so its around 11inch^3 short

OpenStudy (anonymous):

i have another question also

OpenStudy (anonymous):

Kenya exchanges $200 for euros (€). Suppose the conversion rate is €1 = $1.321. How many euros should Kenya receive? €264.20 €151.40 €132.10 €66.05

hartnn (hartnn):

yes, ask :)

OpenStudy (anonymous):

i dont even know what to do

hartnn (hartnn):

€1 = $1.321 so, €2 = $ ?? can u tell ?

OpenStudy (anonymous):

no sir

hartnn (hartnn):

think ? lets say 1 chocolate is worth 4$ what will 2 chocolates worth ?

OpenStudy (anonymous):

8

hartnn (hartnn):

correct!

hartnn (hartnn):

so, here we have €1 is worth $1.321 so, €2 will worth =...?

OpenStudy (anonymous):

2.642?

hartnn (hartnn):

correct! so, €2 is worth $2.642 how much is €200 worth ?

OpenStudy (anonymous):

400

hartnn (hartnn):

no! see, €1 = $1.321 we need $200 worth, lets say its 'x' so, x*1.321 = 200 can you find x from here ?

hartnn (hartnn):

what is 200/1.321 ?

OpenStudy (anonymous):

no :(

OpenStudy (anonymous):

151

OpenStudy (anonymous):

?

hartnn (hartnn):

yes! so its 151.40 :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

i need more help

OpenStudy (anonymous):

sorry for the bother

hartnn (hartnn):

welcome ^_^ you can close this post and ask in another post

hartnn (hartnn):

so that if i can't help you there will be surely someone else to help..